OpenStudy (anonymous):

Christy drove 300 miles on her vacation. She drove an average of 1.25 times faster on the second 150 miles of her trip. Which expression represents the time she spent driving? Let x= her speed on the first half of the trip. a.375/x b. 337.5/x c.270x d. 270/x

OpenStudy (anonymous):

I'm not sure, I wanna say A or B, but like I said I'm not sure.

OpenStudy (anonymous):

I think it would be either d or a though

OpenStudy (anonymous):

im not sure I do online schooling so don't judge little rust at the math stuff

OpenStudy (anonymous):

time = distance/speed. So determine the time she took for the first half = 150/x Second half: 150/1.25x Totaltime taken: 150/x + 150/1.25x = 1/x(150+(150/1.25))= 270/x

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

No prb